OpenStudy (anonymous):

Mary is 14 years younger than her sister Jane. In four years, Jane will be 3 times older than Mary. How old is Jane now? 3 7 17 21

OpenStudy (anonymous):

21

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Just a note, M=J−14 J+4=3(M+4)

OpenStudy (anonymous):

It's actually just systems of equations "complexified" into a word problem.